you have prepared a drug using 2g in 250 ml 0.9% Ns solution for your patient with an irregular heart rhythm. the physician has ordered 2mg/min IV in a continuous infusion via infusion pump. Calculate the ml/hour flow rate. DA equation:
Answer #1

Required volume of medicine = required dose \times available volume \div available dose

                                               = 2 mg \times 250 ml \div 2000 mg

                                               = 0.25 ml

In a solution of 0.25 ml contains 2mg of drug.

Now the infusion pump should be set to 0.25 ml per minute. In one hour the flow rate should be 15 ml (0.25 \times 60)